### v3

# Efficient Multi-agent Navigation with Lightweight DRL Policy

Abstract
This article presents an end-to-end collision avoidance policy based on deep reinforcement learning (DRL) for multi-agent systems, demonstrating encouraging outcomes in real-world applications. In particular, our policy calculates the control commands of the agent based on the raw LiDAR observations. In addition, the number of parameters of the proposed basic model is 140,000, and the size of the parameter file is 3.5 MB, which allows the robot to calculate the actions from the CPU alone. We propose a multi-agent training platform based on a physics-enhanced simulator to further bridge the gap between simulation and the real world. The policy is trained on a policy-gradients-based RL algorithm in a dense and messy training environment. A novel reward function is introduced to address the issue of agents choosing suboptimal actions in some common scenarios. Although the data used for training is exclusively from the simulation platform, the policy can be successfully transferred and deployed in real-world robots. Finally, our policy effectively responds to intentional obstructions and avoids collisions. The website is aviliable in https://sites.google.com/view/xingrong2024efficient/%E9%A6%96%E9%A1%B5
Note to Practitioners
We’ve developed a navigation system that enables robots to navigate collision-free in dynamic environments such as warehouses, factories, and hospitals. It leverages LiDAR data to make rapid decisions with minimal computational resources, making it highly efficient for small robots. This system is trained in physics-enhanced simulations, ensuring its suitability for real-world applications and exceptional performance in handling challenging scenarios like tight spaces and unexpected obstacles. Whether you need robots to move efficiently in a bustling warehouse, transport materials in a factory, or assist in a hospital, this approach provides a practical solution for deploying effective navigation without requiring high-end hardware or intricate configurations.
Index Terms:
Deep reinforcement learning, distributed collision avoidance, reactive and sensor-based planning, simulation to real gap.
I INTRODUCTION
Multi-agent systems are increasingly used in robotics and artificial intelligence, such as industrial automation [1], material transportation [2]. The main challenge is implementing an efficient, robust, and safe navigation algorithm for multiple robots so that the robot can safely and efficiently complete point-to-point movement.
The traditional approaches for solving the problem of multi-agent navigation are primarily categorized into two types. One of the approaches is the centralized approach. These approaches calculate trajectories for each agent through a central server using the state of all agents and their respective workspaces [3]. However, the computational cost and complexity of these approaches increase significantly with the number of agents, and the maintenance of a more extensive communication network. In addition, once the central server is attacked, all agents will stop working. The other approach is the decentralized approach, which is deployed to each agent, empowering them with the capability to make decisions, such as Reciprocal Velocity Obstacle (RVO) [4] and Optimal Reciprocal Collision Avoidance (ORCA) [5]. However, these algorithms rely on an idealized perception of the environment and the states of the other agents, posing challenges for their application in larger-scale scenarios and deployment onto real-world robots. Numerous variants have been proposed to address these limitations [6], [7]. Additionally, RVO and its variants often necessitate tuning numerous hyperparameters to achieve optimal performance, and the adjustment of these parameters is highly dependent on the specific scenario.
Learning-based approaches have been proposed for robot motion planning in recent years [8]. Some approaches use imitation learning to learn from expert data and work successfully [9]. However, the effectiveness of imitation learning depends on expert data quality, and it is challenging to handle scenarios out of the expert data. Other approaches employ reinforcement learning (RL) or deep reinforcement learning (DRL) to train policy. Some RL-based approaches achieve desired outcomes by modifying the reward function [10], [11]. These approaches can yield excellent results in corresponding scenarios, but some suffer from redundant reward function designs and struggle to handle unforeseen scenarios. Some approaches train agents in specific scenarios to learn more diverse policies [12], [13]. Learning in more specific scenarios can enrich policies but demands higher standards for scenario design. Besides, there are efforts to bridge the gap between simulation and the real world by using progressive networks or adding randomness to the simulator [14], [15]. Such approaches enable better deployment of policies learned from simulators on real robots, yet they do not fully leverage physics-enhanced simulators. Finally, as the robot in a decentralized system can only perceive local observations during navigation, the decentralized robot navigation can be defined as the partially observable Markov decision process (POMDP). Most policies use convolutional neural networks (CNN) to encode perceptual information [16], [17], which results in the loss of historical information during navigation. Some approaches propose using long short-term memory (LSTM) [18] to record memory, providing more comprehensive information for navigation [19], [20].
This paper presents a lightweight, end-to-end distributed multi-agent navigation framework that directly maps raw LiDAR observations to agent action commands. LiDAR is selected as the primary perception modality due to its minimal simulation-to-reality gap and compact measurement dimensions, which facilitate effective simulation-to-real transfer [21]. We develop a physics-enhanced simulation platform replicating real-world dynamics to bridge the reality gap in multi-agent training. Extensive experimental validation demonstrates the policy’s robustness in simulated and physical environments, including single-agent and multi-agent scenarios. Notably, the framework successfully handles adversarial conditions, such as rapidly moving human-controlled obstacles.
Contribution. Our main contributions are as follows.
  • We develop a physics-enhanced simulation platform that bridges the simulation-to-reality gap. Comparisons show the physics-integrated system improves environmental randomness and training effectiveness.
  • We propose a lightweight distributed framework using LSTM hidden states to address POMDP information loss. An integrated attention mechanism enhances coordination, with policies trained in stochastic environments and transferable to real-world robots.
  • We introduce a reward function addressing classical formulation limitations, resolving suboptimal behaviors in narrow passages and bidirectional multi-agent navigation.
The rest of this article is organized as follows. Section II is a brief review of the related work. Section III introduces the notation and formulation used in this article, the proposed reward function, the policy’s structure, the overall framework of the approach, and the physics-enhanced simulator platform. Section IV presents the policy’s performance in the simulation and real-world experiments. Finally, we conclude our work and propose directions for future improvement in Section V.
II Related Work
In this section, we briefly describe some related works of multi-agent navigation. These works can be categorized into trajectory-based, reaction-based, and learning-based approaches. Finally, we discuss the approaches to bridge the gap between simulation and the real world.
II-A Trajectory-based Approaches
Trajectory-based approaches generally require detailed environmental information to plan a feasible trajectory for the robot [22]. In recent years, new theories, such as safe corridors, buffered Voronoi cells (BVC), and Graphs of Convex Sets (GCS), have been introduced to transform the navigation problem into a convex optimization problem. Park [23] utilizes a dynamic linear safe corridor while introducing mode-based subgoal planning to tackle planning problems in complex environments. Zhu [24] constructs a chance-constrained buffered uncertainty-aware Voronoi cell (B-UAVC) for each robot and ensures the safety of robot navigation by specifying thresholds. Marcucci [25] proposes a framework that enables convex optimization to effectively and reliably plan trajectories around obstacles through a practical convex relaxation programming problem by using GCS. Although the trajectory-based algorithm can obtain the optimal trajectory, it will lead to significant computational costs or conplex convex optimization when applied in large or unstructured scenarios.
II-B Reaction-based Approaches
Reaction-based approaches accomplish navigation tasks by continuously sensing and responding to the surrounding environment, enabling them to adapt to dynamically changing conditions. Reciprocal velocity obstacles (RVO) [4] maps the dynamic environment into the robot velocity space and generates safe control velocities from these velocity constraints. The ORCA [5] follows the principles of RVO while resolving the jitter problem in RVO and improves obstacle avoidance efficiency. Moreover, ORCA-NH [6] further guarantees smooth, collision-free motions under non-holonomic constraints. However, these approaches struggle to be implemented in the real world because they require a perfect perception of the agent’s surroundings. Still, ORCA and its variants are sensitive to parameters, and the corresponding parameters are different in different scenarios, which makes these approaches challenging to adjust in practical applications.
II-C Reinforcement Learning-based Approaches
RL can be applied well to navigation tasks since the robot navigation problem is a POMDP problem. RL-based multi-agent navigation approaches learn from the interaction, and the quality of the policy depends on the design of the reward function, the interaction quality, and the neural network’s structure.
Generally, the reward function of the navigation problem is similar to the potential field function. Chen [26] proposes a dense reward function that can guide reinforcement learning concisely and effectively. Some approaches define the reward function through other rules. Xie [10] introduces a velocity obstacle term into the reward function, enabling mobile robots to navigate autonomously in spaces filled with static obstacles and dense pedestrian traffic. Some approaches train the policy in specific international scenarios.
Regarding the neural network structure, some policies use CNNs to encode perceptual information [16]. To improve the policy’s ability to leverage historical states, some approaches utilize the hidden state architecture of RNNs to encode and store past information. Liu [19] presents the Decentralized Structured Recurrent Neural Network (DS-RNN), which uses spatial and temporal relationships for robotic decision-making in swarm navigation. Xue [13] introduces a novel multi-agent recurrent deterministic policy gradient algorithm and leverages LSTM to improve action value predictions using historical exploration information.
Some traditional control approaches have also been employed to enhance the safety of RL training. Cheng [27] utilizes a model-based controller leveraging Control Barrier Functions (CBFs) and online learning of unknown system dynamics to ensure safety during the learning phase.
II-D Simulation-to-Real Gap
As training is inefficient in the real world and easily damages the robot, RL-based approaches are usually trained in simulators. However, the gap between simulations and real-world environments remains a crucial barrier preventing the widespread adoption of RL-based approaches in real robots. Several methods have been proposed to address this issue and bridge the simulation-to-reality gap. Carcia [28] proposes an approach to evaluate Lidar sensors and navigation and obstacle avoidance capabilities in simulated scenarios using a real drone platform. Fan [12] proposes a hybrid control RL approach, which initially learns an end-to-end general obstacle avoidance policy through RL. Subsequently, hybrid control is incorporated into this policy to improve its feasibility and robustness in real-world environments.
Our approach further bridges the gap between simulation and the real-world environment by designing a training platform based on a physics-based simulator. We propose a novel policy network based on LSTM and attention mechanism, a state-of-the-art (SOTA) navigation based on a pure LiDAR sensor. Furthermore, a novel reward function is designed to solve common scenarios, such as an agent passing through narrow gaps.
III Method
In this section, we will introduce the notation of this work, the proposed reward function, the architecture of the policy network, the training process, and the simulation platform with training scenarios.
III-A Notations and Problem Statement
This work models the multi-agent navigation process as POMDP. The POMDP can be defined as $<\mathcal{I},\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{R},\mathcal{T},done,\gamma>$, where =$\mathcal{I}=\{1,...,N\}$ denotes $N$ agents, $\mathcal{S}$ denotes the set of states of all agents, $\mathcal{O}=\{\mathcal{O}_{1},...,\mathcal{O}_{N}\}$ denotes the set of observable state of agents, $\mathcal{A}$ denotes the set of actions, $\mathcal{R}$ denotes the reward function, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ denotes the transfer function and $done$ indicates whether an episode has terminated in time or collided with an obstacle, and $\gamma\in[0,1)$ denotes the discount factor. At each time step $t$, the agent’s observation is $o^{t}=[o_{s}^{t},o_{g}^{t},o_{v}^{t}]$, where $o_{s}^{t}\in\mathbb{R}^{5\times L_{n}}$ is formed by concatenating 5 consecutive frames of LiDAR data, with $L_{n}$ being the measurement dimension of the LiDAR data; $o_{g}^{t}\in\mathbb{R}^{2}$ consists of the distance from the agent’s current position to the goal point, and the angle between the agent’s orientation and the direction towards the goal point; $o_{v}^{t}=[v_{i}^{t},\omega_{i}^{t}]\in\mathbb{R}^{2}$ consists of the agent’s linear and angular velocities. The agent $i$ receives $o_{i}^{t}$ and executes action $a_{i}^{t}$ according to policy $\pi_{\theta}(a^{t}_{i}|o^{t}_{i})$, where $\theta$ is the policy parameter. The policy $\pi_{\theta}(a^{t}_{i}|o^{t}_{i})$ is a Gaussian distribution, with the mean calculated by a policy network. The action $a^{t}_{i}=[v^{t}_{i},\omega^{t}_{i}]$ is sampled from the policy $\pi_{\theta}(a^{t}_{i}|s^{t}_{i})$. The state $s_{t}$ takes the action set ${a_{i}^{t},...,a_{N}^{t}}$ to transfer to next state $s_{i}^{t+1}$ according to $\mathcal{T}(\mathcal{S}^{t+1}|\mathcal{S}^{t},a_{i}^{t},...,a_{N}^{t})$. Also, the next observable state set is updated as $\mathcal{T}(O^{t+1}|\mathcal{O}^{t},a_{i}^{t},...,a_{N}^{t})$, and the reward set is obtained by =$\mathcal{R}=\{r(s_{1}^{t},a_{1}^{t}),...,r(s_{N}^{t},a_{N}^{t})\}$. A trajectory of agent $i$ can be represented as $<o_{i}^{1},a_{i}^{1},r_{i}^{1},o_{i}^{2},a_{i}^{2},r_{i}^{2},...,s_{i}^{T},a_{i}^{T},r_{i}^{T}>$, where $T$ represents the number of interaction steps. Each agent $i$ aims to maximize the discounted future reward: $G_{i}^{T}=\sum_{t=0}^{T}\gamma^{t}r(s_{i}^{t},a_{i}^{t})$. The state-value function is defined as $V^{\pi}(s_{i}^{T})=\mathbb{E}[G_{i}^{T}|s^{1}_{i}]$, and the action-value function is given by $Q^{\pi}(s_{i}^{T},a_{i}^{T})=\mathbb{E}[G_{i}^{T}+\gamma V^{\pi}(s_{i}^{T})|s^{1}_{i},a^{1}_{i}]$. An action is chosen as $a={\rm arg\ max}_{a}Q(s,a)$.
III-B Simulation Platform
To bridge the gap between simulation and the real world, we propose a multi-agent training platform based on a physics-enhanced simulator such as Pybullet [29] and Gazebo. In most 2-D simulators, the actions of the agent are usually perfectly executed, and their positions are usually calculated by an equation like $x^{t+1}=x^{t}+v^{t}\Delta t+\frac{1}{2}a^{t}\Delta t^{2}$. In a physics-enhanced simulator, the agent state is computed by the physics engine, i.e., $x^{t+1}=\mathcal{T}(x^{t},v^{t},a^{t},f^{t},...)$, where $\mathcal{T}$ is the transfer function of the physics engine, $f^{t}$ is the force acting on the agent. The transfer function $\mathcal{T}$ makes the execution of the action randomness. A straightforward example is when the straight action executes in different types of simulators, the results differ. the 2-D simulator executes actions without deviation, whereas the agent in the physics-enhanced simulator diverges from the straight path, with even more pronounced deviations observed in real-world robots. This suggests that training in a physics-enhanced simulator enables the agent to learn more realistic policies, thereby improving its adaptability and facilitating seamless deployment on actual robots. To better train the policy, we add Gaussian noise to the agent’s perception to increase randomness further.
(a) 2-D simulator
(b) 3-D simulator
(c) Real robot
Figure 1: Comparison between the simulators and the real world. Agents on all platforms execute commands with $v=1$ m/s, $\omega=0$ rad/s. The direct travel distance of the agent in the simulator is 10 m, and the distance of the real robot is 5 m.
We propose a local replay mechanism for the trajectory collection of the agent’s collision, aiming to speed up the collection of multi-agent interaction trajectories. When the agent collides, the agent will be reset back to the state before the $N$ simulation steps so that the agent can quickly accumulate the collision data and speed up the training, where $N=300$ in this article. We propose this mechanism to address: (1) Spatiotemporal Sparsity: collisions in dense multi-agent scenarios are rare, making global reset methods inefficient; (2) Emergent Obstacle Prediction: sudden disturbances induce conservative behaviors (e.g., slowing/stopping), revealing implicit dynamic obstacle modeling; and (3) Asymmetric Adaptation: resetting single agents forces rapid adaptation to dynamic obstacles, simulating real-world pedestrian unpredictability and enhancing generalization.
Besides, to make the simulator generalize, we adopt the architecture of Gym [30], and we have diligently maintained the interfaces intending to provide users with a more convenient experience when utilizing the simulator. The simulator supports the random addition of random or queued agents or obstacles and controls the movement of each agent through an action set $\mathcal{A}$. At the end of a simulator step, the simulator returns an observed state set $\mathcal{O}$. The update frequency of the simulator can be modified to meet the policy training at different control frequencies. All of these functionalities are designed for easy reconfiguration. For example, to accommodate users seeking to evaluate Safe Reinforcement Learning (Safe-RL) [31], we have reconfigured our custom-designed simulator to include a constrained version by changing two functions of the origin version.
Figure 2: The overall framework diagram of our approaches is presented. The approaches can be divided into trajectory collection and policy update. During trajectory collection, each agent collects the trajectory independently and stores it in a replay buffer. After an action is executed, the simulation platform will return each agent’s reward and observation. After a specific number of iterations, the trajectories stored in the replay buffer will be used to calculate the loss, and the optimizer will then update the network based on the loss through gradient descent. The two stages alternate until a specific number of episodes is reached or the policy converges.
Algorithm 1 PPO For Multi-Agent Training
1:Initialize policy network $\pi_{\theta}$, value function network $V_{\phi}$, and set hyper-parameters
2:for $iteration=1,2,\ldots$ do
3:     for $t=1,2,\ldots,T_{\text{max}}$ do
4:         for each robot $i=1,2,\ldots,N$ do
5:              Run policy network $\pi_{\theta}$ to collect $\{o_{i}^{t},a_{i}^{t},r_{i}^{t}\}$
6:              Run value function network to collect $\{V_{\phi}(o_{i}^{t})\}$
7:         end for
8:     end for
9:     for each robot $i=1,2,\ldots,N$ do
10:         for $t=T_{\text{max}},T_{\text{max}}-1,\ldots,1$ do
11:              Calculate TD-error: $\delta_{i}^{t}=r_{i}^{t}+\gamma V_{\phi}(o_{i}^{t+1})-V_{\phi}(o_{i}^{t})$
12:              Estimate advantages using GAE: $A_{i}^{t}=\sum_{k=t}^{T_{\text{max}}}(\gamma\lambda)^{k-t}\delta_{i}^{k}$
13:              Calculate $V_{i}^{t}=A_{i}^{t}+V_{\phi}(o_{i}^{t})$
14:         end for
15:     end for
16:     Merge and shuffle the collected trajectories
17:     $\theta_{\text{old}}=\theta$, $\phi_{\text{old}}=\phi$
18:     for each epoch in Epoch do
19:         Calculate policy loss:
$L^{P}=-\mathbb{E}\left[\sum_{i=1}^{N}\sum_{t=1}^{T_{\text{max}}}\min\left(\frac{\pi_{\theta}(a_{i}^{t}|o_{i}^{t})}{\pi_{\theta_{\text{old}}}(a_{i}^{t}|o_{i}^{t})}A_{i}^{t},\text{clip}\left(\frac{\pi_{\theta}(a_{i}^{t}|o_{i}^{t})}{\pi_{\theta_{\text{old}}}(a_{i}^{t}|o_{i}^{t})},1-\epsilon,1+\epsilon\right)A_{i}^{t}\right)\right]$
20:         Calculate value loss:
$L^{V}=\max\left(\mathbb{E}\left[\sum_{i=1}^{N}\sum_{t=1}^{T_{\text{max}}}\left((V_{\phi}(o_{i}^{t})-V_{i}^{t})^{2},\left(\text{clip}(V_{\phi}(o_{i}^{t}),V_{\phi_{\text{old}}}(o_{i}^{t})-\epsilon,V_{\phi_{\text{old}}}(o_{i}^{t})+\epsilon)-V_{i}^{t}\right)^{2}\right)\right]\right)$
21:         Calculate entropy loss: $L^{\text{Entropy}}=\mathbb{E}[H(\pi_{\theta}(a_{i}^{t}|o_{i}^{t}))]$
22:         Calculate total loss: $L^{\text{Total}}=L^{P}+\alpha L^{V}+\beta L^{\text{Entropy}}$
23:         Update the parameters of $\pi_{\theta}$ and $V_{\phi}$ using Adam Optimizer with respect to $L^{\text{Total}}$
24:     end for
25:end for
III-C Training Process
We treat distributed multi-agent navigation as plenty of single agents navigating in an unknown environment, meaning each agent collects the trajectory respectively. As shown in Fig. 2, the training process can be divided into trajectory collection and policy update. During the trajectory collection, each agent first observes the LiDAR $o_{s}^{t}$, target point $o_{g}^{t}$, and velocity $o_{v}^{t}$. After that, it utilizes the policy and value network to calculate an action mean $[v^{t},\omega^{t}]$ and a value $v^{t}$. From these, an action $a^{t}$ is sampled based on the policy distribution $\pi_{\theta}(a^{t}|o^{t})$ and executed. The simulator then responds with a reward $r^{t}$, the next observable state $o^{t+1}$, and a termination signal $done^{t}$. The tuple of $<o^{t},a^{t},r^{t},o^{t+1},done^{t},V_{\phi}(o^{t})>$ is stored in replay buffer. This process is repeated until a specified number of iterations have been completed. After the trajectories are collected, the process proceeds to the policy update stage, where the trajectories stored in the replay buffer are used to calculate the total loss $L^{Total}$ for updating the policy and value networks. After updating the networks, the replay buffer will be cleared, and the agents collect trajectories again. The two stages alternate until convergence or episode limit is reached.
Algorithm 1 shows a more detailed training process. In lines 3-9, each agent collects the trajectories respectively and calculates the values using the value network under the same environment. After specific steps, in lines 10-15, according to the trajectories collected by each agent, the corresponding advantages $A$ of each trajectory are calculated through General Advantage Estimation (GAE) [32]. Then, we merge and shuffle the trajectories to get training data. In lines 17-24, we calculate the loss by:
$\begin{array}[]{l}L^{{P}}=-\text{min}(\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{{old}}}(a|s)}A,\text{clip}(\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{{old}}}(a|s)},1-\epsilon,1+\epsilon)A),\\ L^{V}=\max((V_{\theta}-V_{targ})^{2},(\operatorname{clip}(V_{\theta},V_{\theta_{old}}-\varepsilon,\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad V_{\theta_{old}}+\varepsilon)-V_{tart})^{2}),\\ L^{{Entropy}}=H(\pi_{\theta}(a^{t}|o^{t})),\end{array}$ (1)
where $V_{targ}=A+V_{\phi}(s)$. We get the total loss by weighted summing $L^{{Total}}=L^{{P}}+\alpha L^{V}+\beta L^{{Entopy}}$, where $\alpha$ and $\beta$ are hyperparameters. We use the Adam Optimizer [33] to update the policy and value network w.r.t $L^{Total}$. The algorithm proceeds to the next iteration until a specified iteration is reached. We also use some tricks to make the training more stable. The tricks include reward scaling [34], mini-batch updating, advantage normalizing, and learning rate decay.
We use two scenarios during training: single-agent and multi-agent. As we clip the distance input to the network, the size of the scenario does not affect the training. Thus, in the single-agent scenario, the initial point and the goal point of navigation are 10 m apart, and the multi-agent scenario’s initial and goal points are 16 m apart. The two scenarios are shown in Fig. 3.
(a) Single-agent scenario
(b) Multi-agent scenario
[…]
The measurement dimension $L_{n}$ of the LiDAR of the policy is different, and the max length of the LiDAR beam is 4 m. We first train the policy with $L_{n}=36$, which works successfully in the single-agent scenario. In the multi-agent scenario, the $L_{n}$ is updated to 72. Because if $L_{n}=36$, the agent may lose information when other agents stay in the internal of the LiDAR beam, with the biggest internal = 0.34 m, while $L_{n}=72$, the biggest internal = 0.17 m is enough for most situations. The field of view (FOV) is set to $\pi$ rad. Additionally, we stack the LiDAR observations of 5 consecutive frames to provide more comprehensive observation information. The agent’s velocity is set within the $[0,1]$ m/s, and the angular velocity is set within the $[-\pi,\pi]$ rad/s. The control frequency used in training is 60 Hz, which can make the reward variation of each action more apparent and maintain efficient control. Finally, the maximum number of simulation steps is set to 2500, sufficient for the agent to finish navigation through the optimal path.
III-D Reward Function
The dense reward function used in RL-based navigation algorithms consists of distance reward, obstacle penalty, and efficiency reward. An example of such a reward function is:
$\begin{array}[]{l}R=r_{g}+r_{c}+r_{other},\\ r_{g}=\left\{\begin{array}[]{ll}r_{{arrival}}\qquad\qquad\quad\ \ \text{if }\left\|p_{i}^{t}-g_{i}\right\|<0.1,\\ w_{g}\left(\left\|p_{i}^{t-1}-g_{i}\right\|-\left\|p_{i}^{t}-g_{i}\right\|\right)\quad\text{otherwise,}\end{array}\right.\\ r_{c}=\left\{\begin{array}[]{ll}r_{{collision}}\qquad\qquad\text{if }d_{i}^{t}<0.01,\\ w_{c}(0.1-d_{i}^{t})\qquad\text{ if }d_{i}^{t}<0.1,\\ 0\qquad\qquad\qquad\quad\text{ otherwise,}\end{array}\right.\\ \end{array}$ (2)
where $r_{other}$ is the reward that varies in different approaches, $r_{g}$ is the distance reward, which is computed as the difference in the distance between the robot and the goal point at the current time step and the previous time step, $r_{c}$ is the obstacle penalty, $w_{g}$ and $w_{c}$ are the weight of $r_{g}$ and $r_{c}$ respectively, $d_{i}^{t}$ means the smallest distance between agent $i$ and obstacles at time step $t$. However, this reward function setup may make agents learn conservative policy. Additionally, many works impose control constraints, such as velocity and angular velocity, on $r_{other}$ to maintain action coherence. Too many constraints may lead to instability in the training process and difficulty adjusting the weights between various terms. Moreover, in some scenarios, the reward function does not adequately indicate optimal actions.
(a) Agent passes through narrow gaps.
(b) Multi-agent traveling in opposite directions.
Figure 4: Scenarios where the classical reward function demonstrates significant limitations.
As shown in Fig. 4, when an agent needs to navigate through the gap between two obstacles, it may incur penalties due to $r_{c}$ when approaching them, even though traversing the gap is the optimal action. As a result, the agent may circumnavigate obstacles. Furthermore, agents focus exclusively on proximal threats in multi-agent scenarios while ignoring potentially hazardous agents at greater distances, resulting in unsafe navigation decisions. To address the issues above, we have introduced modifications to the $r_{c}$ component, specifically by:
$\begin{array}[]{l}r^{\prime}_{c}=\left\{\begin{array}[]{ll}r_{collision}&\text{if }d_{i}^{t}<0.01,\\ w^{\prime}_{c}\left(l-sl_{i}^{t}\right)+r_{c}&\text{ otherwise,}\end{array}\right.\end{array}$ (3)
where $r_{arrival}=5$, $r_{collision}=-5$, $w_{g}=5$, $w^{\prime}_{c}=-0.25$, $w_{c}=-0.03$, $l$ represents the maximum length of the LiDAR rays, $sl_{i}^{t}$ represents the LiDAR observations of agent $i$ at time step $t$, $w^{\prime}_{c}$ is a weight vector that varies with angular velocity, with the weight decreasing in $\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-\omega\Delta t)^{2}}{2\sigma^{2}}}$ fashion from $\omega\Delta t$ towards both sides, where $\sigma=0.2$, $\Delta t$ is the time interval of the control. The schematic diagram of $w^{\prime}_{c}$ is shown in Fig. 5.
Figure 5: The weight of the scan penalty of the $r^{\prime}_{c}$.
Using $r^{\prime}_{c}$, the policy pays more attention to the direction of the current angular velocity while maintaining the closest distance to the surrounding obstacles. The reward based on angular velocity of the agent allows the policy to better cope with the mentioned situations and enable the agent to pay more attention to the choice of angular velocity.
III-E Network Structure
The agent only requires more precise, goal-oriented actions when nearing the target location, whereas general directional movement suffices otherwise. We clip the agent-goal distance rather than feeding raw distance values into the network to enable the policy to differentiate between these two regimes explicitly. This design decouples the policy’s decision-making from absolute distance metrics, ensuring scalability to larger environments while maintaining consistent behavioral logic across spatial scales.
To address the loss of information in POMDP, we first employ LSTM Layers to encode LiDAR observations. By stacking sequential LiDAR observations, the LSTM module captures historical observation memory and integrates consecutive observation patterns. While the policy performs well in static environments and 6-agent test scenarios, its performance degrades significantly when the number of agents exceeds 6. This limitation arises because the LSTM-based policy overly focuses on LiDAR observations while neglecting the agent’s velocity, making it difficult to navigate effectively in denser multi-agent environments.
To address this issue, we introduce an attention mechanism that dynamically balances the importance of LiDAR observations and velocity information. First, the velocity and goal information are encoded through fully connected (FC) layers to produce the query vector $Q$, while the LiDAR observations are processed by the LSTM module with FC layers to generate the value $V$ and key $K$ vectors, respectively. The attention mechanism then integrates all relevant information into a comprehensive representation, as described in (4). Integrating the proposed reward function enhances policy and achieves robust navigation in complex multi-agent scenarios. The architecture, which combines LSTM with an attention mechanism, is designated the LSTM-Attention policy.
TABLE I: Experiment with 1 robot and 15 obstacles.
success rate(%) collision rate(%) trap rate (%) average step
IJRR [12] 91.3 8.6 0.1 1383.23
  Linear policy 67.2 2.4 30.4 550.40
LSTM policy 93.1 6.9 0.0 1187.73
LSTM-Attention 95.9 3.9 0.2 1300.78
  (Safe) LSTM-Attention 97.6 2 0.04 1353.63
TABLE II: Experiment with 6 robots and 5 obstacles.
success rate(%) collision rate(%) trap rate (%) average step
NH-ORCA [6] 98.4 0.33 1.27 1198.82
IJRR [12] 85.03 14.63 0.33 1617.84
LSTM policy 91.89 8.15 0.0 981.01
LSTM-Attention 94.78 4.68 0.53 1805.94
As shown in Fig. 2, we use a three-layer LSTM module to encode the laser measurements $o_{s}^{t}$ with both hidden sizes = 256 and get LSTM encoding $S$. Then, $K$ and $V$ are computed using two FC layers, each with the output sizes = 256, applied to the encoding $S$. At the same time, $[o_{g}^{t},\ o_{v}^{t}]$ passes through two FC layers, which output sizes = 256, to obtain $Q$. Then, the encoding of the LiDAR observations, $X$, is obtained by multi-head attention:
$\begin{array}[]{l}\text{MultiHead}(Q,K,V)=\text{Concat}(\text{head}_{1},\ldots,\text{head}_{h})W^{O},\\ \text{where }\text{head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}),\\ \text{Attention}(Q,K,V)=\text{Softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V,\end{array}$ (4)
where $d_{k}$ is the size of $K$. The final encoding is obtained by the weighted sum of $X$ and $S$ with an adaptive residual adjustment, $\alpha X+(1-\alpha)S$, where $\alpha$ updates with the network. In actor and critic, the final encoding is decoded by two FC layers with output size = 128 and output size = 2 to get the mean of the policy $[v^{t},\omega^{t}]$, and another two FC layers with output size = 128 and output size = 1 to get the value network $V_{\phi}(o^{t})$. Note that each LSTM layer is followed by a rectified linear unit (ReLU), and so is the FC layer. Additionally, the last FC layer of the actor and critic modules has no ReLU layers.
IV EXPERIMENT EVALUATION
IV-A Experimental Setup
All models are implemented with PYTORCH on a server with a GTX4090 GPU and CPU of Intel Xeon 8336C. We test the model and select their optimal parameter settings.
To evaluate the ability of the policy, we test it in both simulation and the real world compared with NH-ORCA [6] and the policy of [12]. In the following section, we refer to the policy of [12] by the CNN policy and the measurement dimension of the LiDAR $L_{n}=512$. The LSTM-Attention policy is the proposed policy. We conduct an ablation experiment by removing the attention mechanism from the proposed policy, which is the LSTM policy. Additionally, we replace the LSTM layer with the Linear layer and obtain the Linear policy to validate the effectiveness of the LSTM layer.
In the simulation tests, we use four metrics to evaluate the policy. The metrics consist of the success rate, the collision rate, the trap rate, and the average step. The success rate means the agent’s success rate navigates to the goal in time without collision. The collision rate means the rate at which the agent collides with other objects. The trap rate means the rate at which agent cannot reach the goal in time and without collision. The average step is the average simulation step of the agent to reach the goal successfully. We use simulation steps instead of navigation cost time as the metric because the latter depends on the computational speed of the simulator, which may not accurately reflect real-world navigation time. Finally, we perform all simulation tests 1000 times.
Figure 6: Multi-agent test without obstacle.
IV-B Simulation Test
[…]
Figure 7: Real-world single-agent experiment. The TurtleBot3 first passes through narrow passages, crosses through the obstacles, and finally reaches the goal point. The TurtleBot3 reaches the goal point in 46 s.
IV-C Real-World Test
To evaluate the policy’s performance in real-world scenarios, we deploy it on a TurtleBot3 platform. The TurtleBot3’s maximum linear and angular velocities are limited to 0.2 m/s and 1 rad/s, respectively; thus, we scale the policy’s output to align with these constraints. Additionally, the control frequency is set to 10 Hz. The LiDAR sensor on the TrutleBot3 provides FOV $=2\pi$ rad with 360 beams, but we only utilize 72 beams within a reduced FOV $=0.8\pi$ rad to match our experimental requirements. We align the rest of the policy with the simulation settings. We use the wheel odometry and LiDAR sensor on the TurtleBot3 for perception. Wheel odometry calculates the robot’s relative position and orientation to the goal, while the LiDAR sensor captures the surrounding environment observations.
In the single-agent scenario, we evaluate the policy’s performance by navigating the TurtleBot3 through narrow passages. The TurtleBot3 successfully passes the passages and avoids obstacles of unknown positions and shapes while reaching the goal point. As illustrated in Fig. 7, the TurtleBot3 begins in a room, traverses a narrow passage to enter the goal room in 22 s, and then avoids obstacles to achieve the goal point in an additional 24 s. Finally, the TurtleBot3 completes the navigation task as it approaches the goal point.
In the multi-agent scenario, we evaluate the policy’s performance in avoiding other agents and responding to sudden disturbances. Each TurtleBot3 needs to navigate through obstacles while avoiding the other robots and human. In this evaluation, the human acts as a dynamic disturbance by intentionally interfering with the TurtleBot3s. As shown in Fig. 8, the human first appears in front of a TurtleBot3 to force it to alter its trajectory, then crosses directly between two TurtleBot3s. Subsequently, the human continues interfering with one TurtleBot3 until it reaches its goal point. Despite these disturbances, all TurtleBot3s successfully avoid collisions with each other and the human.
The navigation algorithm demonstrates robust performance in both single-agent and multi-agent scenarios. In the single-agent scenario, the TurtleBot3 successfully navigates through narrow passages and avoids obstacles of unknown positions and shapes, efficiently reaching the goal point. In multi-agent scenario, the algorithm enables TurtleBot3s to avoid collisions with other robots and respond to dynamic disturbances, such as intentional human interference. The TurtleBot3s dynamically adjust their trajectories to accommodate sudden obstacles while maintaining collision-free navigation. These results highlight the algorithm’s adaptability and reliability in handling static and dynamic challenges, ensuring safe and efficient navigation in real-world applications.
Figure 8: Real-world multi-agent experiment. The red pattern indicates the state of TurtleBot3, the blue pattern indicates the state of a human, the arrow represents the direction of velocity, and the circle represents reaching the goal point (TurtleBot3) or remaining stationary (human). At t = 0 s, each TurtleBot3 navigates to the goal point. A human appears in the environment at t = 5 s, starting to interfere with TurtleBot3. The human first forces a TurtleBot3 to change its trajectory and crosses in front of two TurtleBot3s at t = 10 s and t = 13 s. Then, the human interferes with another TurtleBot3 again until it reaches the goal point for the rest of the time. All TurtleBot3s reach the goal points in 31 s.
V CONCLUSION AND FUTURE WORK
We present a lightweight, end-to-end multi-agent navigation framework that directly maps raw LiDAR observations to action commands, eliminating the need for explicit obstacle classification while enabling efficient navigation through constrained spaces. Our approach leverages a physics-enhanced simulation platform to bridge the simulation-to-real gap, facilitating practical multi-agent training. The reinforcement learning-based solution demonstrates robust performance across simulated and real-world environments, including challenging scenarios with adversarial obstructions, while maintaining computational efficiency.
Our work makes several key contributions to the field of multi-agent navigation. We develop a distributed navigation policy capable of operating in stochastic environments, incorporating an LSTM-based architecture to address information loss in POMDPs. The proposed physics-integrated training platform enables safe and efficient RL validation, complemented by a novel reward function that effectively resolves suboptimal action selection in complex scenarios. Comprehensive evaluations in diverse environments demonstrate the framework’s effectiveness, robustness, and real-world applicability.
In the future, we will identify multiple promising directions for future research. These include extending the framework to handle dynamic traffic scenarios and unmanned aerial vehicle (UAV) applications, refining the reward function to address complex edge cases such as U-shaped obstacles, and enhancing policy robustness in adversarial environments. The proposed framework’s adaptability and performance suggest strong potential for broader applications in autonomous navigation systems.